### v8

Abstract
Constructing efficient, tightly secure short signature schemes based on lattices has remained a significant challenge in cryptographic theory. In their ASIACRYPT 2016 paper, Boyen and Li raised an open problem: how to construct a tightly or almost-tightly secure homomorphic signature scheme in the standard model under the Short Integer Solution (SIS) hardness assumption. Under the same assumption, this paper presents the first lattice-based short signature scheme that achieves adaptive security and provides a proof of its almost-tight security. Furthermore, the proposed scheme retains the potential to be further refined into a tightly secure signature scheme.
keywords:
lattice, linearly homomorphic signature, tightly secure, standard model, SIS problem
1 Introduction
Homomorphic signatures refer to a type of signature that allows any entity to perform homomorphic operations on authenticated data, enabling the generation of new data. The entity can then obtain a valid signature for the new data without using the signing private key. Since homomorphic signatures were put forward, they have attracted more and more attention. The unique properties of homomorphic signatures provide a vast theoretical research space and significant scientific research value, playing an important role in many practical applications. For example, they can offer effective solutions in fields such as network coding[2, 3, 4, 35, 36] and cloud computing[5, 6, 37, 38, 39].
The concept of homomorphic signature schemes was first introduced by R. Rivest [7] in a presentation at Cambridge University in 2000. Shortly thereafter, in 2002, Johnson et al. [8] provided the formal definition and security analysis of homomorphic signatures. Since then, various types of homomorphic signature schemes have emerged. When classifying homomorphic signature schemes based on the homomorphic operation functions, they can be categorized into linear homomorphic signatures, polynomial function homomorphic signatures, and fully homomorphic signatures.
In 2007, Zhao et al. proposed the first linearly homomorphic signature scheme [2], which allows arbitrary linear combinations of signed data, facilitating the verification of message integrity and effectively preventing contamination attacks on applications based on network coding. In 2009, Boneh et al. [9] provided the formal definition and security model for linearly homomorphic signature schemes and constructed a linearly homomorphic signature scheme for packet authentication in network coding protocols based on bilinear groups. Subsequently, further improvements were made in terms of efficiency, security, and privacy protection, leading to the proposal of numerous efficient and practical linearly homomorphic signature schemes. Early research primarily based linearly homomorphic signatures on number-theoretic assumptions, such as the discrete logarithm problem or the RSA problem [9, 10, 11, 12, 13, 40, 41, 42]. These schemes are efficient but must be designed over $\mathbb{F}_{p}$ (where $p$ is a large prime) to ensure the difficulty of the RSA or discrete logarithm problem, and they cannot resist quantum computing attacks.
To resist quantum attacks, Boneh and Freeman[14] first proposed a linearly homomorphic signature scheme over $\mathbb{F}_{2}$ (the binary field) at the PKC 2011 conference, with its security based on the $k$-SIS problem ($k$-small integer solution problem). Subsequently, at the EUROCRYPT 2011 conference, Boneh and Freeman[15] introduced another linearly homomorphic signature scheme, whose security relies on the hardness of the SIS problem, and proposed a homomorphic signature scheme for polynomial functions, with security based on ideal lattices. In 2013, Wang et al.[16] improved the first linearly homomorphic signature scheme based on lattices, resulting in a new linearly homomorphic signature whose security is directly based on the standard SIS problem. All these schemes[14, 15, 16] have been proven secure in the random oracle model.
However, in reality, the ideal random oracle model is difficult to achieve. Chen et al. combined bonsai tree techniques[17] with the intersection method of two integer lattices proposed in [14], and for the first time constructed a lattice-based linearly homomorphic signature scheme over small fields in the standard model[18]. Building on Chen’s work, Lin et al. constructed two lattice-based linearly homomorphic signature schemes in the standard model by utilizing full-rank differential hash functions and linearly homomorphic chameleon hash functions[19].
In the security reduction of a cryptographic system, if the loss in the adversary’s advantage is only a constant factor, the system is said to be tightly secure; if the security loss is a logarithmic factor of the security parameter, it is referred to as almost tightly secure(as discussed in Section 4). Tight security is a highly desirable property in the design of cryptographic schemes, as it provides stronger practical security guarantees. Regarding this security property, Boyen and Li [20] posed an open problem at the 2016 Asiacrypt conference (ASIACRYPT 2016): How to construct a short, homomorphic signature scheme with tight or almost tight security in the standard model, based on the hardness of the Short Integer Solution (SIS) problem. In 2024, Chen et al. [21] designed the first lattice-based almost tightly secure linearly homomorphic signature scheme by building upon the scheme in literature [18], incorporating pseudorandom function techniques[20] and the key-homomorphic algorithm Eval [22, 23]. However, the scheme fails to achieve almost tight security under the existential unforgeability against chosen-message attacks (EUF-CMA) security model(see Definition 3 for details). Under the standard model, constructing a homomorphic signature scheme with tight security based on the SIS problem remains an open problem, particularly in achieving tight or almost security under the standard EUF-CMA security model.
1.1 Our Contribution
This study is dedicated to addressing the aforementioned open problem. Specifically, we propose a novel lattice-based linearly homomorphic signature scheme. The proposed scheme is rigorously proven to satisfy the existential unforgeability against chosen-message attacks (EUF-CMA) security in the standard model, with its security reduction achieving almost tightness. In addition, the framework of our proposed scheme holds the potential for further optimization to achieve full tight security.
1.2 Our Techniques
Reference [19] proposed a signature scheme in the standard model based on the SIS problem, featuring adaptive security, short signature size, and almost tight security; it further indicated that a tightly secure short signature scheme can be constructed given a tightly secure pseudorandom function (PRF). Since the scheme proposed in this paper incorporates it as a component of the entire signature scheme, it is necessary to present a concise summary of its technical key points. Specifically, the scheme first generates a series of random matrices: a left matrix $\boldsymbol{A}$, subspace selection matrices $\boldsymbol{A}_{0}$ and $\boldsymbol{A}_{1}$, a set of matrices $\{\boldsymbol{B}_{i}\}$ corresponding to the PRF key, and message-representation matrices $\boldsymbol{C}_{0}$ and $\boldsymbol{C}_{1}$, with the PRF implemented as a NAND circuit. The private key comprises a short basis $\mathbf{T}_{\boldsymbol{A}}$ for $\boldsymbol{A}$ and the PRF key $K$. During signing, a matrix $\boldsymbol{A}_{\text{PRF},M}$ corresponding to the PRF circuit is computed via key homomorphism, which is then combined with the message to obtain the matrix $\boldsymbol{F}_{M,1-b}$; using the short basis $\mathbf{T}_{\boldsymbol{A}}$, a low-norm vector $\mathbf{d}_{M}$ satisfying $\boldsymbol{F}_{M,1-b}\cdot\mathbf{d}_{M}\equiv\mathbf{0}\pmod{q}$ is generated as the signature. Verification checks that $\mathbf{d}_{M}$ is non-zero and has sufficiently small norm, and confirms whether $\boldsymbol{F}_{M,b}\cdot\mathbf{d}_{M}\equiv\mathbf{0}\pmod{q}$ or $\boldsymbol{F}_{M,1-b}\cdot\mathbf{d}_{M}\equiv\mathbf{0}\pmod{q}$ holds—if all conditions are satisfied, the signature is accepted.
Before detailing the signature generation process, this paper first introduces a newly proposed core technique, the decomposition algorithm denoted as $\textbf{Decompose}(\mathbf{m})$. This algorithm decomposes the majority of vectors $\mathbf{m}\in\mathbb{F}_{2}^{2n}$ into a pair of vectors $(\mathbf{u},\mathbf{v})$ with specific Hamming weights, where the Hamming weight of $\mathbf{u}$ is fixed at n1n-1, and that of $\mathbf{v}$ is either $n$ or $n+1$, such that $\mathbf{x}=\mathbf{u}+\mathbf{v}$. The pivotal role of this algorithm lies in the security proof, ensuring that the simulated signatures are statistically indistinguishable from the genuine signatures. This guarantees that the adversary cannot differentiate between the simulated environment and the real environment during forgery attacks. Given that the related analysis involves intricate statistical properties, the detailed discussion will be presented in Theorem 8 (Security Proof).
Furthermore, it should be specifically noted that the signature in the proposed scheme consists of two components, namely $(\sigma_{\tau},\sigma)$. Here, $\sigma_{\tau}$ can be viewed as a signature generated for the tag $\tau$ based on the scheme from Reference [19], serving to resist tampering and forgery attacks targeting the tag, while $\sigma$ is the signature generated for the message itself. When performing homomorphic operations on signatures, only the $\sigma$ component is manipulated, with the $\sigma_{\tau}$ component remaining constant. For instance, if the signatures corresponding to messages $m_{1}$ and $m_{2}$ are $(\sigma_{\tau},\sigma_{1})$ and $(\sigma_{\tau},\sigma_{2})$, respectively, then the signature for the combined message $m_{1}+m_{2}$ can be obtained through homomorphic evaluation: $(\sigma_{\tau},\sigma_{1})+(\sigma_{\tau},\sigma_{2})=(\sigma_{\tau},\sigma_{1}+\sigma_{2})$.
Finally, we elaborate on the signature generation process in detail. Our signature scheme begins by generating a matrix $\mathbf{A}$ along with its trapdoor $\mathbf{T_{A}}$ (serving as the private key) via the trapdoor generation algorithm $\textbf{TrapGen}(q,h,n)$. Concurrently, $k$ (where $k$ is an even number) linearly independent vectors $\{\mathbf{\alpha}_{1},\dots,\mathbf{\alpha}_{k}\}$ are selected from $\mathbb{Z}_{q}^{h}$. These vectors, together with $\mathbf{A}$ and the public key from Reference [19], constitute the public key of our scheme. For a message $\mathbf{m}_{i}\in\mathbb{F}_{2}^{k}$ to be signed and its corresponding tag $\tau\in\{0,1\}^{n}$, the signature algorithm from Reference [19] is first employed to obtain a signature on the tag $\tau$. Subsequently, the algorithm $\text{Decompose}(m_{i})$ is invoked to decompose the message $m_{i}$ into a pair of vectors $(\mathbf{u}_{i},\mathbf{v}_{i})$. Provided that both $\mathbf{u}_{i}$ and $\mathbf{v}_{i}$ are non-zero vectors, the linear combinations $t(\mathbf{u}_{i})=\sum_{j=1}^{k}u_{ij}\mathbf{\alpha}_{j}$ and $t(\mathbf{v}_{i})=\sum_{j=1}^{k}v_{ij}\mathbf{\alpha}_{j}$ are computed. The preimage sampling algorithms $\textbf{SamplePre}(\mathbf{A},\mathbf{T}_{\mathbf{A}},t(\mathbf{u}_{i}),V)$ and $\textbf{SamplePre}(\mathbf{A},\mathbf{T}_{\mathbf{A}},t(\mathbf{v}_{i}),V)$ are then called to generate the signature components $\mathbf{\sigma}(\mathbf{u}_{i})$ and $\mathbf{\sigma}(\mathbf{v}_{i})$, respectively, where $V$ is the Gaussian parameter governing the sampling distribution. Finally, the signature is output as $\mathbf{\sigma}_{i}=\mathbf{\sigma}(\mathbf{u}_{i})+\mathbf{\sigma}(\mathbf{v}_{i})$. Finally, the signature is output as $\boldsymbol{\sigma}_{i}=\boldsymbol{\sigma}(\mathbf{u}_{i})+\boldsymbol{\sigma}(\mathbf{v}_{i})$.
2 Preliminaries
2.1 Notation
We use the notations $\mathcal{O}$, $\widetilde{\mathcal{O}}$, and $\omega$ to characterize the growth of functions. Let $f$ and $g$ be two functions of $n$. We say that $f(n)=\mathcal{O}(g(n))$ if and only if there exist a constant $c$ and an integer $N$ such that $f(n)\leq cg(n)$ for all integers $n>N$. Similarly, $f(n)=\widetilde{\mathcal{O}}(g(n))$ holds if and only if $f(n)=\mathcal{O}(g(n)\cdot\log^{c^{\prime}}n)$ for some constant $c^{\prime}$. $f(n)=\omega(g(n))$ if and only if there exists an integer $N$ such that $g(n)\leq cf(n)$ for any constant $c>0$ and any integer $n>N$. Let $n$ denote the security parameter. If $f(n)=\mathcal{O}(n^{c})$ for a constant $c>0$, then we denote $f(n)=\mathrm{poly}(n)$. If for all $c>0$, the function $f(n)$ is $\mathcal{O}(n^{-c})$, then this function is called negligible and denoted as $\mathrm{negl}(n)$. If the probability of an event occurring is $1-\mathrm{negl}(n)$, we say that the event holds with overwhelming probability.
The capital letter $\mathbb{Z}$ denotes the ring of integers, $\mathbb{Z}_{q}$ denotes the ring of integers modulo $q(q\geq 2)$ and $\mathbb{R}^{n}$ denotes the $n$-dimensional Euclidean space. We then let capital letters $A,B,C$, etc. represent matrices, and bold lowercase letters $\mathbf{a},\mathbf{b},\mathbf{c}$ etc. represent vectors. If $A=(\mathbf{a}_{1},...,\mathbf{a}_{n})\in\mathbb{R}^{h\times n}$, the norm of the matrix $A$ is defined as $\parallel A\parallel=\max_{1\leq i\leq n}\parallel\mathbf{a}_{i}\parallel$, where $\parallel\mathbf{a}_{i}\parallel$ represents the $l_{2}$ norm of the vector. We further denote $\widetilde{A}=(\widetilde{\mathbf{a}}_{1},...,\widetilde{\mathbf{a}}_{n})$ as the Gram-Schmidt orthogonalization of $A$, that is,
$\widetilde{\mathbf{a}}_{1}=\mathbf{a}_{1},\quad\widetilde{\mathbf{a}}_{i}=\mathbf{a}_{i}-\sum_{j=1}^{i-1}\frac{\langle\mathbf{a}_{i},\widetilde{\mathbf{a}}_{j}\rangle}{\langle\widetilde{\mathbf{a}}_{j},\widetilde{\mathbf{a}}_{j}\rangle}\widetilde{\mathbf{a}}_{j},2\leq i\leq n,$
where $\langle,\rangle$ represents the standard inner product in Euclidean space.
For any distribution $\mathcal{D}$, $x\sim\mathcal{D}$ means that $x$ follows the distribution $\mathcal{D}$, and $x\leftarrow\mathcal{D}$ means that the sampling of a random value by the distribution. For a set $\mathcal{X}$, let $x\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\mathcal{X}$ denote uniformly randomly choosing $x$ form $\mathcal{X}$. For any probabilistic polynomial-time (PPT) algorithm $Alg$, we write $y\leftarrow Alg(x)$ to denote that the algorithm takes input $x$ and then outputs the result $y$.
2.2 Linearly homomorphic signature: definition and security model
Before introducing the linearly homomorphic signature, we define a multi-data homomorphic signature [21, 26, 27]. Different from the classical definition of digital signatures, this homomorphic signature consists of five polynomial-time algorithms (HSetup, HKeyGen, HSign, HVerify, Evaluate), allowing the signer to sign multiple different data sets (each data set corresponds to a tag), and the size of each data set can be different. It is also assumed that the verifier knows the tag of the data set over which they want to verify the computation. The definition of the multi-data homomorphic signature is as follows.
Definition 1.
(Multi-data homomorphic signature, [21, 26, 27]) A multi-data homomorphic signature scheme is a tuple of probabilistic, polynomial-time algorithms (HSetup, HKeyGen, HSign, Verify, Evaluate) as follows:
  1. $\bullet$
[…]
  2. $\bullet$
    HKeyGen(pp)($p$). Taking the public parameter $pp$ as input, this algorithm outputs a pair of public key and private key $(pk,sk)$
  3. $\bullet$
[…]
  4. $\bullet$
[…]
  5. $\bullet$
[…]
  1. (1)
    For all tags $\tau\in\{0,1\}^{n}$, all $\mathbf{m}\in\mathcal{M}$, and all $i\in\{1,...,k\}$, if $\sigma\longleftarrow\textbf{HSign}(sk,\tau,\mathbf{m},i)$, then
    $\textbf{HVerify}(pk,\tau,\mathbf{m},\sigma,\pi_{i})=1.$
  2. (2)
    For all tags $\tau\in\{0,1\}^{n}$, all $\overrightarrow{\mathbf{m}}=(\mathbf{m}_{1},...,\mathbf{m}_{k})\in\mathcal{M}^{k}$, and all $f\in\mathcal{F}$, if $\sigma_{i}\longleftarrow\textbf{HSign}(sk,\tau,\mathbf{m},i)$, for   $i\in\{1,...,k\}$, then
    $\textbf{HVerify}(pk,\tau,f(\overrightarrow{\mathbf{m}}),\textbf{Evaluate}(pk,\tau,f,(\sigma_{1},...,\sigma_{k})))=1.$
We say that a signature scheme as above is multi-data $\mathcal{F}$-homomorphic.
Remark 1: Typically, we require the message dataset $\mathbf{m}_{1},...,\mathbf{m}_{k}$ to be a set of linearly independent vectors.
The definition of linearly homomorphic signatures is given below.
Definition 2.
([21]) In the above definition of a multi-data homomorphic signature scheme, if the message space $\mathcal{M}$ consists of $n$-dimensional vectors over a ring $\mathbf{R}$, and the set of admissible functions $\mathcal{F}$ consists of $\mathbf{R}$-linear function from $(\mathbf{R^{n}})^{k}$ to $\mathbf{R}$, we refer to it as the multi-data linearly homomorphic signature scheme. A linear function $f(\mathbf{v}_{1},...,\mathbf{v}_{k})=\sum_{i=1}^{i=k}c_{i}\mathbf{v}_{i}$ is represented as the vector $(c_{1},...,c_{k})\in\mathbf{R}^{k}$.
Referring to [15, 19], the security model of linearly homomorphic signature is given below. (Note: The signature scheme definition provided in [15, 19] is essentially no different from Definition 2)
Definition 3.
(EUF-CMA, [15, 19]) A linearly homomorphic signature $\mathcal{LS}$=(HSetup,HKeyGen,HSign,HVerify, Evaluate) satisfies existential unforgeability under chosen-message attacks (EUF-CMA) if for all $k$ the advantage of any probabilistic, polynomial-time adversary $\mathcal{A}$ in the following game is negligible in the security parameter $n$:
HSetup: The challenger runs $\textbf{Setup}(1^{n},k)$ to obtain $(pk,sk)$ and gives $pk$ to $\mathcal{A}$. The public key defines a message space $\mathcal{M}$, A signature space $\Sigma$, and a set $\mathcal{F}$ of linear function $f:\mathcal{M}^{k}\longrightarrow\mathcal{M}$.
Queries: Proceeding adaptively, $\mathcal{A}$ specifies a sequence of data sets $\overrightarrow{\mathbf{m}}_{i}\in\mathcal{M}^{k}$. For each $i$, the challenger choose $\tau_{i}$ uniformly from $\{0,1\}^{n}$ and gives to $\mathcal{A}$ the tag $\tau_{i}$ and the signature $\sigma_{i,j}$ for j=1,,kj=1,...,k.
Output: $\mathcal{A}$ output a tag $\tau^{*}\in\{0,1\}^{n}$, a message $\mathbf{m}^{*}\in\mathcal{M}$, a linear function $f\in\mathcal{F}$, and a signature $\sigma^{*}\in\Sigma$.
The adversary wins if $\textbf{HVerify}(pk,\tau^{*},\mathbf{m}^{*},\sigma^{*},f)=1$ and either
  1. (1)
    $\tau^{*}\neq\tau_{i}$ for all $i$ (I-type forgery), or
  2. (2)
    $\tau^{*}=\tau_{i}$ for some $i$ but $\mathbf{m}^{*}\neq f(\overrightarrow{\mathbf{m}}_{i})$ (II-type forgery).
The advantage of $\mathcal{A}$ is the probability that $\mathcal{A}$ wins the security game.
In reference [19], the authors also defined a weaker security model, namely existential unforgeability against non-adaptive chosen-message attacks (EUF-naCMA). This model requires the adversary A to submit the set of messages to be signed to the challenger before the attack begins. Due to space limitations and the fact that this model falls outside the scope of this paper, its detailed definition will not be elaborated here. Readers may refer to [19] for further details.
In [21], the purpose of Chen is to construct a almost tightly secure linearly homomorphic signature Scheme. However, in practice, it cannot satisfy almost tight security under EUF-CMA. Therefore, inspired by [28, 29], they proposed unforgeability under the selective-tag static chosen-message-attack (U-ST-SCMA), and achieved tight security under this security model.
Definition 4.
(U-ST-SCMA, [21]) For a linearly homomorphic signature scheme $\mathcal{LS}$=(HSetup,HKeyGen,HSign, HVerify, Evaluate), we consider the following game:
Initialization: (select tag and data phase): Let $n_{s}$ be the number of the data sets . The challenger uniformly samples $(\tau_{i}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,1\}^{n})_{i\in\{1,...,n_{s}\}}$ and sends them to the adversary $\mathcal{A}$. Then, $\mathcal{A}$ specifies a sequence of data sets $(\overrightarrow{\mathbf{m}}_{i}=(\mathbf{m}_{i1},...,\mathbf{m}_{ik})\in\mathcal{M}^{k})_{i\in\{1,...,n_{s}\}}$ and a target tag $\tau_{g}\in\{\tau_{1},...,\tau_{n_{s}}\}$, and returns them.
HSetup: The challenger runs $\textbf{HSetup}(1^{n},k)$ to obtain $(pk,sk)$ and gives $pk$ to $\mathcal{A}$.
Queries: The challenger computes $(\sigma_{ij})_{i\in\{1,...,n_{s}\},j\in\{1,...,k\}}$, and gives the signatures $(\sigma_{ij})_{i\in\{1,...,n_{s}\},j\in\{1,...,k\}}$ to $\mathcal{A}$.
Output: $\mathcal{A}$ outputs a tag $\tau^{*}\in\{0,1\}^{n}$, a message $\mathbf{m}^{*}\in\mathcal{M}$, a linear function $f\in\mathcal{F}$, and a signature $\sigma^{*}\in\Sigma$. Then the adversary wins if $\textbf{HVerify}(pk,\tau^{*},\mathbf{m}^{*},\sigma^{*},f)=1$ and either
  1. (1)
    $\tau^{*}\neq\tau_{i}$ for all $i$ (I-type forgery), or
  2. (2)
    $\tau^{*}=\tau_{g}$ but $\mathbf{m}^{*}\neq f(\overrightarrow{\mathbf{m}}_{g})$ (II-type forgery).
Compared to EUF-CMA, U-ST-SCMA is a weaker security form, as this scheme requires the pre-disclosure of the messages to be signed and the target tags to be forged before providing the public key to the adversary.
In homomorphic signatures, privacy also needs to be considered. Below, we provide the definition of weakly context hiding[15].
Definition 5.
(Weakly context hiding, [15]) A linearly homomorphic signature scheme $\mathcal{LS}$=(HSetup,HKeyGen,HSign,HVerify, Evaluate) is weakly context hiding if for all $k$, the advantage of any probabilistic, polynomial-time adversary $\mathcal{A}$ in the following game is negligible in the security parameter $n$:
HSetup: The challenger runs $\textbf{HSetup}(1^{n},k)$ to obtain $(pk,sk)$ and gives $pk$ and $sk$ to $\mathcal{A}$. The public key defines a message space $\mathcal{M}$, a signature space $\Sigma$, and a set $\mathcal{F}$ of linear function $f:\mathcal{M}^{k}\longrightarrow\mathcal{M}$.
Challenge: $\mathcal{A}$ outputs $(\overrightarrow{\mathbf{m}}_{0},\overrightarrow{\mathbf{m}}_{1},f_{1},...,f_{s})$ with $\overrightarrow{\mathbf{m}}_{0},\overrightarrow{\mathbf{m}}_{1}\in\mathcal{M}^{k}$. The function $f_{1},...,f_{s}\in\mathcal{F}$ and satisfy
$f_{i}(\overrightarrow{\mathbf{m}}_{0})=f_{i}(\overrightarrow{\mathbf{m}}_{1})\quad\text{for all}\quad i=1,...,s.$
In response, the challenger generates a random bit $b\in\{0,1\}$ and a random tag $\tau\in\{0,1\}^{n}$. It signs the message in $\overrightarrow{\mathbf{m}}_{b}$ using the tag $\tau$ to obtain a vector $\overrightarrow{\sigma}$ of $k$ signatures. Next, for i=1,,si=1,...,s the challenger computes a signature $\sigma_{i}=\textbf{Evaluate}(pk,\tau,f_{i},\overrightarrow{\sigma})$ of $f_{i}(\overrightarrow{\mathbf{m}}_{b})$. It sends the $(\tau,\sigma_{1},...,\sigma_{s})$ to $\mathcal{A}$. Note that the functions $f_{1},...,f_{s}$ can be output adaptively after $\overrightarrow{\mathbf{m}}_{0}$, $\overrightarrow{\mathbf{m}}_{1}$ are output.
Output: $\mathcal{A}$ output a bit $b^{\prime}$. The adversary $\mathcal{A}$ wins the game if $b^{\prime}=b$. The advantage of $\mathcal{A}$ is the probability that $\mathcal{A}$ wins the game.
We say that the linearly homomorphic signature scheme is $s$-weakly context hiding if the attacker cannot win the privacy game after seeing at most $s$ signatures derived from $\overrightarrow{\mathbf{m}}_{0}$, $\overrightarrow{\mathbf{m}}_{1}$.
2.3 Entropy and Statistical Distance
Definition 6.
(Statistical distance, [30]) Let $M\subset\mathbb{R}^{n}$ be a finite or countable set, and let $X$ and $Y$ be two discrete random variables taking values in $M$. The statistical distance between $X$ and $Y$ is defined as:
$\bigtriangleup(X,Y)=\frac{1}{2}\sum_{a\in M}|P\{X=a\}-P\{Y=a\}|.$
When the statistical distance between two distributions is less than a negligible value, we say that the two distributions are statistically close.
Definition 7.
([26], Min-entropy) For a random variable $X$, its min-entropy is defined as:
$H_{\infty}(X)=-\log\left(\max_{x\in X}\Pr[X=x]\right).$
The average conditional min-entropy of a random variable $X$ conditional on a correlated variable $Y$ is defined as:
$H_{\infty}(X|Y)=-\log\left(\mathbb{E}_{y\in Y}\left\{\max_{x\in X}\Pr[X=x|Y=y]\right\}\right).$
The optimal probability of an unbounded adversary guessing $X$ given the correlated value $Y$ is $2^{-H_{\infty}(X|Y)}$[26].
Lemma 1.
([26]) Let $X$, $Y$ be arbitrarily random variables where the support of $Y$ lies in $\mathcal{Y}$, Then
$H_{\infty}(X|Y)>H_{\infty}(X)-\log(|\mathcal{Y}|).$
2.4 Background on Lattices and Hard Problems
Definition 8.
(Lattice, [30] Let $\Lambda\subset\mathbb{R}^{n}$ be a non-empty subset. $\Lambda$ is called a lattice if:
(1) it is an additive subgroup of $\mathbb{R}^{n}$;
(2) there exists a positive constant $\lambda=\lambda(\Lambda)>0$ such that
$\mathrm{min}\{\parallel\mathbf{x}\parallel|\mathbf{x}\in\Lambda,\mathbf{x}\neq 0\}=\lambda.$
$\lambda$ is called the minimum distance.
A full-rank $n$-dimensional lattice can also be expressed as a linear combination of a set of basis vectors $B=\{\mathbf{b}_{1},...,\mathbf{b}_{n}\}\subset\mathbb{R}^{n}$:
$\Lambda=\mathcal{L}(B)=\{B\cdot\mathbf{x}=\sum_{i=1}^{n}x_{i}\mathbf{b}_{i}|x=(x_{1},...,x_{n})^{\top}\in\mathbb{Z}^{n}\}.$
We call $\Lambda^{*}$ the dual lattice of $\Lambda$ if
$\Lambda^{*}=\left\{\mathbf{y}\in\mathbb{R}^{n}\mid\langle\mathbf{y},\mathbf{x}\rangle\in\mathbb{Z},\quad\text{for all }\mathbf{x}\in\Lambda\right\}.$
Definition 9.
( $q$-ary lattices, [15]) Let $A\in\mathbb{Z}_{q}^{n\times m}$, $u\in\mathbb{Z}^{n}$. The following two $q$-ary lattices are defined as:
[…]
The set $\Lambda_{q}^{\mathbf{u}}$ is a coset of $\Lambda_{q}^{\bot}$ since $\Lambda_{q}^{\mathbf{u}}=\Lambda_{q}^{\bot}+\mathbf{t}$ for any $\mathbf{t}$ such that $A\cdot\mathbf{t}\equiv\mathbf{u}(\mathrm{mod}q)$.
Definition 10.
(Short integer solution, [31]) Let $n$, $m$, $q$ be positive integers, with $m=\mathrm{poly(}n)$. Let $A\in\mathbb{Z}_{q}^{n\times m}$ be a uniformly distributed random matrix over $\mathbb{Z}_{q}$, and let $\beta\in\mathbb{R}$ such that $0<\beta<q$. The SIS problem is to find a short integer solution $\mathbf{x}$ satisfying the following condition:
$A\cdot\mathbf{x}\equiv 0(\mathrm{mod}q),\quad\mathrm{and}\quad\mathbf{x}\neq 0,\parallel\mathbf{x}\parallel\leq\beta.$
We denote the SIS problem with parameters $\textbf{SIS}_{q,n,m,\beta}$ or $\textbf{SIS}_{q,\beta}$.
Theorem 1.
(Worst-case to average-case reduction, [32]) In $\mathbb{Z}_{q}^{n\times m}$, $m=poly(n)$, $q\geq\beta\omega(\sqrt{n\log n})$, and $\beta=poly(n)$, if there is a polynomial time algorithm that solves $\textbf{SIS}_{q,\beta}$ with non-negligible probability, then there is a polynomial time algorithm that finds short non-zero vectors, which are only a $\lambda\geq\widetilde{\mathcal{O}}(\sqrt{n})$ factor longer than the shortest vector, in all lattices of dimension $n$.
2.5 Discrete Gaussian Distributions, Lattice Trapdoor Algorithms, and Sampling Algorithms
Definition 11.
(Discrete Gaussian Distribution, [30]) Let $s$ be a positive real number and $\mathbf{c}\in\mathbb{R}^{n}$ be a vector. The Gaussian function centered at $\mathbf{c}$ with parameter $s$ is defined as: $\rho_{s,\mathbf{c}}(\mathbf{x})=e^{-\frac{\pi}{s^{2}}\parallel\mathbf{x}-\mathbf{c}\parallel^{2}}.$ The discrete Gaussian measure $\mathcal{D}_{\Lambda,s,\mathbf{c}}$ defined on the lattice $\Lambda$ is given by:
$\mathcal{D}_{\Lambda,s,\mathbf{c}}(\mathbf{x})=\frac{\rho_{s,\mathbf{c}}(\mathbf{x})}{\rho_{s,\mathbf{c}}(\Lambda)},$
where $\rho_{s,\mathbf{c}}(\Lambda)=\sum_{\mathbf{x}\in\Lambda}\rho_{s,\mathbf{c}}(x)$.
Micciancio and Goldwasser [33] proved that a full-rank set $S$ in the lattice $\Lambda$ can be transformed into a basis $T$ such that both have similarly low Gram-Schmidt norms.
Lemma 2.
([33], Lemma 7.1) Let $\Lambda$ be an $n$-dimensional lattice. There is a deterministic polynomial-time algorithm that, given an arbitrary basis of and a full-rank set $S=\{s_{1},...,s_{n}\}$ in $\Lambda$, returns a basis $T$ of $\Lambda$ satisfying
$\parallel\widetilde{T}\parallel\leq\parallel\widetilde{S}\parallel\text{and}\parallel T\parallel\leq\parallel S\parallel\frac{\sqrt{n}}{2}.$
Theorem 2.
([25],Theorem 3.1) Let $q\geq 3$ be odd, $h$ be a positive integer, and let $n:=\lceil 6h\log q\rceil.$ There is a probabilistic polynomial-time algorithm TrapGen(q,h,n)(q,h,n) that outputs a pair $(A\in\mathbb{Z}_{q}^{h\times n},T_{A}\in\mathbb{Z}^{n\times n})$ such that $A$ is statistically close to a uniform rank $h$ matrix in $\mathbb{Z}_{q}^{h\times n}$, and $T_{A}$ is a basis for $\Lambda_{q}^{\bot}(A)$ satisfying
$\parallel\widetilde{T_{A}}\parallel\leq\mathcal{O}(\sqrt{h\log q})\quad\text{and}\quad\parallel T_{A}\parallel\leq\mathcal{O}(h\log q).$
Lemma 3.
(Sampling from Discrete Gaussian Distribution, [32]) Let $q\geq 2$ , $A\in\mathbb{Z}_{q}^{n\times m}$ and let $T_{A}$ be basis for $\Lambda_{q}^{\bot}(A)$ and $s\geq\widetilde{T_{A}}\cdot\omega(\sqrt{\log m})$. Then for $\mathbf{c}\in\mathbb{R}^{n}$ and $\mathbf{u}\in\mathbb{Z}^{n}$:
  1. 1.
[…]
  2. 2.
[…]
When $s\geq\omega(\sqrt{\log n})$, we denote the Gaussian sampling algorithm over the integer lattice $\mathbb{Z}^{n}$ as $\textbf{SampleDom}(1^{n},s)$. That is, when $\mathbf{x}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\textbf{SampleDom}(1^{n},s)$, $\mathbf{x}$ is statistically close to the distribution $\mathcal{D}_{\mathbb{Z}^{n},s}$. And these preimages have a conditional min-entropy of $\omega(\log n)$[32].
Lemma 4.
([32]) Let $n$ and $h$ be positive integers, and $q$ be a prime, such that $n\geq 2h\lg q$. Then for all but a $2q^{-h}$ fraction of all $A\in\mathbb{Z}_{q}^{h\times n}$ and for any $s\geq\omega(\sqrt{\log n})$, the distribution of the $\alpha=A\cdot\mathbf{x}(\mathrm{mod}q)$ is statistically close to uniform over $\mathbb{Z}_{q}^{h}$, where $\mathbf{x}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\textbf{SampleDom}(1^{n},s)$.
Lemma 5.
([34]) Let $\Lambda$ be an $n$-dimensional lattice, and $T$ be a basis of the lattice $\Lambda$. If $s\geq\parallel\widetilde{T}\parallel\cdot\omega(\sqrt{\log n})$, then for any $\mathbf{c}\in\mathbb{R}^{n}$, we have:
$\mathrm{Pr}\{\parallel\mathbf{x}-\mathbf{c}\parallel>s\sqrt{n}:\mathbf{x}{\leftarrow}\mathcal{D}_{\Lambda,s,\mathbf{c}}\}\leq\mathrm{negl}(n).$
Definition 12.
( Smoothing parameter[34]) For any $n$-dimensional lattice $\Lambda$ and any given $\epsilon>0$, the smoothing parameter of the lattice is defined as
$\eta_{\epsilon}(\Lambda)=\min\left\{s>0\mid\rho_{\frac{1}{s}}(\Lambda^{*})<1+\epsilon\right\}.$
Lemma 6.
([32]) Let $q\geq 3$, $h$ and $n$ be positive integers satisfying $n\geq 2h\lg q$. Then there exists a negligible function $\epsilon(n)$ such that $\eta_{\epsilon}(\Lambda_{q}^{\bot}(A))<\omega(\sqrt{\log n})$ for all but at most a $q^{-h}$ fraction of $A$ in the $\mathbb{Z}_{q}^{h\times n}$ .
Lemma 7.
([43]) For any $n$-dimensional lattice $\Lambda$, $\mathbf{c}\in\text{span}(\Lambda)$, real number $\epsilon\in(0,1)$, and $s\geq\eta_{\epsilon}(\Lambda)$, it holds that
$\underset{\mathbf{x}\sim D_{\Lambda,s,\mathbf{c}}}{\Pr}\left[\|\mathbf{x}-\mathbf{c}\|>s\sqrt{n}\right]\leq\frac{1+\epsilon}{1-\epsilon}\cdot 2^{-n}$
2.6 Statistical Properties Related to Discrete Gaussian Distributions
Theorem 3.
([14]) Let $\Lambda$ be an $n$-dimensional lattice, $A\in\mathbb{Z}^{s\times k}$, and $T\in\mathbb{Z}^{k\times n}$. $X$ is a random matrix of size $k\times n$ where the rows are independent, and the $i$-th row follows the distribution $\mathcal{D}_{\Lambda+\mathbf{t}_{i},\sigma}$, with $\mathbf{t}_{i}$ being the $i$-th row of the matrix $T$. If $\sigma>\mu(A)\eta_{\epsilon}(\Lambda)$ where $\epsilon$ is some negligible value and $\mu(A)=\sqrt{\mathrm{det}(AA^{\top})}$. Then the random variable $AX$ is statistically close to a distribution parameterized by $\Lambda$, $\sigma$, $A$, ATAT.
This theorem shows that for any matrix $A\in\mathbb{Z}^{s\times k}$ and sufficiently large $\sigma$, the random variable $AX$ is statistically close to a distribution parameterized by $\Lambda,\sigma,A$ ,ATAT. Furthermore, the dependence on the matrix $T$ is only reflected through the quantity $AT$. Therefore, if $X_{1}$ and $X_{2}$ are random variables defined in the same way as $X$, but using matrices $T_{1}$ and $T_{2}$, then when $AT_{1}=AT_{2}$, $AX_{1}$ and $AX_{2}$ are statistically close.
Lemma 8.
([14]) $\mathbf{t}_{i}\in\mathbb{Z}^{m}$ and $\mathbf{x}_{i}$ are mutually independent random variables sampled from a Gaussian distribution $D_{\mathbf{t}_{i}+\Lambda,\sigma}$ over $\mathbf{t}_{i}+\Lambda$ for $i=1,2,\cdots,k$ in which $\Lambda$ is a lattice and $\sigma\in\mathbb{R}$ is a parameter. Let $\mathbf{c}=(c_{1},\cdots,c_{k})\in\mathbb{Z}^{k}$ and $g=\mathrm{gcd}(c_{1},\cdots,c_{k})$, $\mathbf{t}=\sum_{i=1}^{k}c_{i}\mathbf{t}_{i}$. If $\sigma>\|\mathbf{c}\|\eta_{\epsilon}(\Lambda)$ for some negligible number $\epsilon$, then $\mathbf{z}=\sum_{i=1}^{k}c_{i}\mathbf{x}_{i}$ statistically closes to $D_{\mathbf{t}+g\Lambda,\|\mathbf{c}\|\sigma}$.
Lemma 9.
Let $A\in\mathbb{Z}_{q}^{n\times m},s>0,\mathbf{u}\in\mathbb{Z}^{m}$, and $\Lambda=\Lambda_{q}^{\mathbf{u}}(A)$. If $\mathbf{x}$ is sampled from $\mathcal{D}_{\mathbb{Z}^{m},s}$ conditioned on $A\mathbf{x}\equiv\mathbf{u}\pmod{q}$, then the conditional distribution of $\mathbf{x}$ is $\mathcal{D}_{\Lambda,s}$.
Proof.
According to Definition 9, the discrete Gaussian distribution defined on $\Lambda$ is:
$\mathcal{D}_{\Lambda,s}(x)=\frac{\rho_{s}(\mathbf{x})}{\rho_{s}(\Lambda)},\quad\mathbf{x}\in\Lambda,\text{ where }\rho_{s}(\mathbf{x})=e^{-\frac{\pi}{s^{2}}\|\mathbf{x}\|^{2}}.$
Then, for $\mathbf{x}\sim\mathcal{D}_{\mathbb{Z}^{m},s}$ and $\mathbf{x}\in\Lambda$, it can be regarded as the following conditional distribution:
$\mathcal{D}_{\mathbb{Z}^{m},s}(\mathbf{x}\mid\mathbf{x}\in\Lambda)=\frac{\mathcal{D}_{\mathbb{Z}^{m},s}(\mathbf{x})}{\mathcal{D}_{\mathbb{Z}^{m},s}(\Lambda)}=\frac{\rho_{s}(x)/\rho_{s}(\mathbb{Z}^{m})}{\rho_{s}(\Lambda)/\rho_{s}(\mathbb{Z}^{m})}=\frac{\rho_{s}(\mathbf{x})}{\rho_{s}(\Lambda)}=\mathcal{D}_{\Lambda,s}(\mathbf{x}).$
Therefore, $\mathbf{x}\sim\mathcal{D}_{\Lambda,s}$.
Theorem 4.
Assume that $s_{1}\geq s_{2}\geq\eta_{\epsilon}(\mathbb{Z}^{n})$, where $\epsilon=\frac{1}{\mathrm{poly}(n)}$ is a sufficiently small value. If $|\frac{1}{s_{1}^{2}}-\frac{1}{s_{2}^{2}}|\leq\frac{\epsilon}{2\pi ns_{1}^{2}}$ and $(\frac{s_{2}}{s_{1}})^{n}\geq 1-\frac{\epsilon}{1-\epsilon}$, then the statistical distance between $D_{\mathbb{Z}^{n},s_{1}}$ and $D_{\mathbb{Z}^{n},s_{2}}$ is at most $3\epsilon$.
Proof.
By the Poisson summation formula, we have:
$\rho_{s}(\mathbb{Z}^{n})=\sum_{\mathbf{x}\in\mathbb{Z}^{n}}\exp\left(-\pi\frac{\|\mathbf{x}\|^{2}}{s^{2}}\right)=s^{n}\sum_{\mathbf{x}\in\mathbb{Z}^{n}}\exp\left(-\pi\|\mathbf{x}\|^{2}s^{2}\right).$
Since $s>\eta_{\epsilon}(\mathbb{Z}^{n})$ and the dual lattice of $\mathbb{Z}^{n}$ is itself, it follows that:
$\rho_{s}(\mathbb{Z}^{n})=s^{n}\sum_{\mathbf{x}\in\mathbb{Z}^{n}}\exp\left(-\pi\|\mathbf{x}\|^{2}s^{2}\right)=s^{n}\rho_{1/s}(\mathbb{Z}^{n})\leq s^{n}(1+\epsilon).$
Thus, we obtain the inequality:
$s_{i}^{n}\leq\rho_{s_{i}}(\mathbb{Z}^{n})\leq s_{i}^{n}(1+\epsilon).$
Consider the ratio $\frac{\rho_{s_{2}}(\mathbb{Z}^{n})}{\rho_{s_{1}}(\mathbb{Z}^{n})}$. Noting that $\frac{s_{2}}{s_{1}}\leq 1$, we have:
$\left(\frac{s_{2}}{s_{1}}\right)^{n}\frac{1}{1+\epsilon}\leq\frac{\rho_{s_{2}}(\mathbb{Z}^{n})}{\rho_{s_{1}}(\mathbb{Z}^{n})}\leq\frac{(1+\epsilon)s_{2}^{n}}{s_{1}^{n}}\leq 1+\epsilon.$
Let $U=\{\mathbf{x}\mid\|\mathbf{x}\|\leq 1\}$ and $B=\{\mathbf{x}\mid\|\mathbf{x}\|\leq s_{1}\sqrt{n}\}$. For sufficiently small $\epsilon$ (e.g., $\epsilon=1/\text{poly}(n)$), by Theorem 6 (for large $n$), we get:
$\Pr_{\mathbf{x}\sim\mathcal{D}_{\mathbb{Z}^{n},s_{1}}}\left[\|\mathbf{x}\|>s_{1}\sqrt{n}\right]=\frac{\rho_{s_{1}}(\mathbb{Z}^{n}\setminus s_{1}\sqrt{n}U)}{\rho_{s_{1}}(\mathbb{Z}^{n})}=\sum_{\mathbf{x}\notin B}D_{\mathbb{Z}^{n},s_{1}}(\mathbf{x})\leq\frac{1+\epsilon}{1-\epsilon}\cdot 2^{-n}\leq\epsilon.$
Similarly,
$\sum_{\mathbf{x}\notin B}D_{\mathbb{Z}^{n},s_{2}}(\mathbf{x})=\frac{\rho_{s_{2}}(\mathbb{Z}^{n}\setminus s_{1}\sqrt{n}U)}{\rho_{s_{2}}(\mathbb{Z}^{n})}\leq\frac{\rho_{s_{2}}(\mathbb{Z}^{n}\setminus s_{2}\sqrt{n}U)}{\rho_{s_{2}}(\mathbb{Z}^{n})}\leq\frac{1+\epsilon}{1-\epsilon}\cdot 2^{-n}\leq\epsilon.$
Thus,
$\sum_{\mathbf{x}\notin B}\left|\mathcal{D}_{\mathbb{Z}^{n},s_{1}}(\mathbf{x})-\mathcal{D}_{\mathbb{Z}^{n},s_{2}}(\mathbf{x})\right|\leq\sum_{\mathbf{x}\notin B}\mathcal{D}_{\mathbb{Z}^{n},s_{1}}(\mathbf{x})+\sum_{\mathbf{x}\notin B}\mathcal{D}_{\mathbb{Z}^{n},s_{2}}(\mathbf{x})\leq 2\epsilon.$
For $\mathbf{x}\in B$, consider the ratio:
$\frac{\rho_{s_{1}}(\mathbf{x})}{\rho_{s_{2}}(\mathbf{x})}=\exp\left(-\pi\|\mathbf{x}\|^{2}\left(\frac{1}{s_{1}^{2}}-\frac{1}{s_{2}^{2}}\right)\right).$
Notice that:
$\left|\pi\|\mathbf{x}\|^{2}\left(\frac{1}{s_{1}^{2}}-\frac{1}{s_{2}^{2}}\right)\right|\leq\pi s_{1}^{2}n\cdot\frac{\epsilon}{2\pi ns_{1}^{2}}=\frac{\epsilon}{2}.$
Since $|e^{z}-1|\leq 2|z|$ for $|z|\leq\frac{1}{2}$, we have:
$1-\epsilon\leq\frac{\rho_{s_{1}}(\mathbf{x})}{\rho_{s_{2}}(\mathbf{x})}\leq 1+\epsilon.$
For $\mathbf{x}\in B$, it follows that:
$\frac{\mathcal{D}_{\mathbb{Z}^{n},s_{1}}(\mathbf{x})}{\mathcal{D}_{\mathbb{Z}^{n},s_{2}}(\mathbf{x})}=\frac{\rho_{s_{1}}(\mathbf{x})}{\rho_{s_{2}}(\mathbf{x})}\cdot\frac{\rho_{s_{2}}(\mathbb{Z}^{n})}{\rho_{s_{1}}(\mathbb{Z}^{n})}\subset\left[\left(\frac{s_{2}}{s_{1}}\right)^{n}\frac{1-\epsilon}{1+\epsilon},(1+\epsilon)^{2}\right]\subset[1-3\epsilon,1+4\epsilon].$
Hence,
$\left|\mathcal{D}_{\mathbb{Z}^{n},s_{1}}(\mathbf{x})-\mathcal{D}_{\mathbb{Z}^{n},s_{2}}(\mathbf{x})\right|\leq 4\epsilon\mathcal{D}_{\mathbb{Z}^{n},s_{1}}(\mathbf{x}),$
and therefore:
$\sum_{\mathbf{x}\in B}\left|\mathcal{D}_{\mathbb{Z}^{n},s_{1}}(\mathbf{x})-\mathcal{D}_{\mathbb{Z}^{n},s_{2}}(\mathbf{x})\right|\leq 4\epsilon\sum_{\mathbf{x}\in B}\mathcal{D}_{\mathbb{Z}^{n},s_{1}}(\mathbf{x})\leq 4\epsilon.$
Combining these results, we calculate:
$\frac{1}{2}\sum_{\mathbf{x}\in\mathbb{Z}^{n}}$ $\left|\mathcal{D}_{\mathbb{Z}^{n},s_{1}}(\mathbf{x})-\mathcal{D}_{\mathbb{Z}^{n},s_{2}}(\mathbf{x})\right|$
$=\frac{1}{2}\sum_{\mathbf{x}\in B}\left|\mathcal{D}_{\mathbb{Z}^{n},s_{1}}(\mathbf{x})-\mathcal{D}_{\mathbb{Z}^{n},s_{2}}(\mathbf{x})\right|$
$\quad+\frac{1}{2}\sum_{\mathbf{x}\notin B}\left|\mathcal{D}_{\mathbb{Z}^{n},s_{1}}(\mathbf{x})-\mathcal{D}_{\mathbb{Z}^{n},s_{2}}(\mathbf{x})\right|$
$\leq\frac{1}{2}(4\epsilon+2\epsilon)$
$\leq 3\epsilon.$
Corollary 5.
Let $k=\mathrm{poly}(n)$ be even, $q=\mathrm{poly}(n)\geq(nk)^{2}$, and $V=\sqrt{2n\log q}\log n\geq\omega(\log n)$. Then $\mathcal{D}_{\mathbb{Z}^{n},V}$ and $\mathcal{D}_{\mathbb{Z}^{n},\sqrt{\frac{k\pm 2}{k}}V}$ are statistically indistinguishable.
Proof.
Let $s_{1}=V$ and $s_{2}=\sqrt{\frac{k-2}{k}}V$. We proceed to show that $\mathcal{D}_{\mathbb{Z}^{n},s_{1}}$ and $\mathcal{D}_{\mathbb{Z}^{n},s_{2}}$ are statistically indistinguishable.
By Lemma 6 and the construction of $V$, there exists a sufficiently small $\epsilon$ such that $s_{1}>\eta_{\epsilon}(\mathbb{Z}^{n})$ and $s_{2}>\eta_{\epsilon}(\mathbb{Z}^{n})$. For convenience, let $\epsilon=\frac{1}{n^{d}}$ where $d>0$ and $k=n^{d+2}$. When $n$ is sufficiently large, we have:
$\left|\frac{1}{s_{1}^{2}}-\frac{1}{s_{2}^{2}}\right|=\left|\frac{1}{V^{2}}-\frac{1}{\frac{k-2}{k}V^{2}}\right|=\frac{2}{(k-2)V^{2}}\leq\frac{1}{2\pi n^{d+1}V^{2}}=\frac{\epsilon}{2\pi ns_{1}^{2}}.$
Next, applying Bernoulli’s inequality: for any $x\in(0,1)$ and $y\geq 1$, $(1-x)^{y}\geq 1-xy$. Thus,
$\left(\frac{s_{2}}{s_{1}}\right)^{n}$ $=\left(1-\frac{2}{k}\right)^{n/2}=\left(1-\frac{2}{n^{d+2}}\right)^{n/2}$ (1)
$\geq 1-\frac{2}{n^{d+2}}\cdot\frac{n}{2}=1-\frac{1}{n^{d+1}}$
$\geq 1-\frac{1}{n^{d}-1}=1-\frac{1}{\frac{1}{\epsilon}-1}=1-\frac{\epsilon}{1-\epsilon}.$
By Theorem 4, $\mathcal{D}_{\mathbb{Z}^{n},V}$ and $\mathcal{D}_{\mathbb{Z}^{n},\sqrt{\frac{k-2}{k}}V}$ are statistically indistinguishable.
Following identical steps, we can show that $\mathcal{D}_{\mathbb{Z}^{n},V}$ and $\mathcal{D}_{\mathbb{Z}^{n},\sqrt{\frac{k+2}{k}}V}$ are also statistically indistinguishable. This completes the proof. ∎
3 New linearly homomorphic signature Scheme
Before introducing the specific construction methods, it is first necessary to propose a decomposition algorithm named Decompose. The detailed content is as follows:
Theorem 6.
Let $k$ be an odd integer with $k=\mathrm{poly}(n)$. Define the sets
$\mathcal{X}$ $=\{x\in\mathbb{F}_{2}^{2k}\mid\|x\|=k-1\}\sqcup\{0\},$
$\mathcal{Y}$ $=\{y\in\mathbb{F}_{2}^{2k}\mid\|y\|=k\}\sqcup\{0\},$
$\mathcal{Z}$ $=\{z\in\mathbb{F}_{2}^{2k}\mid\|z\|=k+1\}\sqcup\{0\}.$
There exists a polynomial-time algorithm $\textbf{Decompose}(\mathbf{w})$ that, on input $\mathbf{w}\in\mathbb{F}_{2}^{2k}$, outputs a pair $(\mathbf{u},\mathbf{v})$ satisfying $\mathbf{w}=\mathbf{u}+\mathbf{v}$ with $\mathbf{u}\in\mathcal{X}$ and $\mathbf{v}\in\mathcal{Y}$ or $\mathcal{Z}$. Furthermore, when $\mathbf{w}\neq 0$, at most one of $\mathbf{u}$ and $\mathbf{v}$ is zero.
Proof.
For any $\mathbf{w}\in\mathbb{F}_{2}^{2k}$, let $\|\mathbf{w}\|=t$. The algorithm proceeds as follows:
Case 1: $t$ is odd
(1) If $t=k$, set $\mathbf{u}=0$ and $\mathbf{v}=\mathbf{w}$. Clearly, $\mathbf{w}=\mathbf{u}+\mathbf{v}$ with $\mathbf{u}\in\mathcal{X}$ and $\mathbf{v}\in\mathcal{Y}$, so the algorithm outputs $(\mathbf{u},\mathbf{v})$.
(2) If $1\leq t<k$, suppose the coordinates of $\mathbf{w}$ equal to 1 are $i_{1},i_{2},\dots,i_{t}$ where $1\leq i_{s}\leq 2k$ for $s=1,\dots,t$.
  • $\bullet$
    Compute a pair $(n,m)$ such that $t=n+m$ and nm=1n-m=1;
  • $\bullet$
    Construct $\mathbf{v}$: Select $n$ coordinates from $i_{1},i_{2},\dots,i_{t}$ and set them to 1. Choose knk-n coordinates from the remaining 2kt2k-t coordinates and set them to 1. All other coordinates are 0. Denote this vector as $\mathbf{v}$;
  • $\bullet$
    Construct $\mathbf{u}$: Set the remaining $m$ coordinates in $i_{1},i_{2},\dots,i_{t}$ (not selected for $\mathbf{v}$) to 1, and set the $n$ selected coordinates to 0. All other coordinates match $\mathbf{v}$. Denote this vector as $\mathbf{u}$.
Verification: Clearly, $\mathbf{w}=\mathbf{u}+\mathbf{v}$. Moreover, $\|\mathbf{u}\|=m+k-n=k-(n-m)=k-1$ and $\|\mathbf{v}\|=n+(k-n)=k$, so $\mathbf{u}\in\mathcal{X}$ and $\mathbf{v}\in\mathcal{Y}$. The algorithm outputs $(\mathbf{u},\mathbf{v})$.
(3) If $k<t\leq 2k-1$, suppose the coordinates of $\mathbf{w}$ equal to 1 are $i_{1},i_{2},\dots,i_{t}$.
  • $\bullet$
    Compute a pair $(n,m)$ such that $t=n+m$ and nm=1n-m=1;
  • $\bullet$
    Construct $\mathbf{v}$: Select $n$ coordinates from $i_{1},i_{2},\dots,i_{t}$ and set them to 1. Choose $\frac{2k-1-t}{2}$ coordinates from the remaining 2kt2k-t coordinates and set them to 1. All other coordinates are 0. Denote this vector as $\mathbf{v}$;
  • $\bullet$
    Construct $\mathbf{u}$: Set the remaining $m$ coordinates in $i_{1},i_{2},\dots,i_{t}$ (not selected for $\mathbf{v}$) to 1, and set the $n$ selected coordinates to 0. All other coordinates match $\mathbf{v}$. Denote this vector as $\mathbf{u}$.
Verification: Clearly, $\mathbf{w}=\mathbf{u}+\mathbf{v}$. Moreover,
$\|\mathbf{u}\|$ $=m+\frac{2k-1-t}{2}=m+\frac{2k-1-(n+m)}{2}=k-1,$
$\|\mathbf{v}\|$ $=n+\frac{2k-1-t}{2}=n+\frac{2k-1-(n+m)}{2}=k,$
so $\mathbf{u}\in\mathcal{X}$ and $\mathbf{v}\in\mathcal{Y}$. The algorithm outputs $(\mathbf{u},\mathbf{v})$.
Case 2: $t$ is even
(1) If $t=0$, set $\mathbf{u}=\mathbf{w}$ and $\mathbf{v}=\mathbf{w}$. This trivially satisfies the output conditions, so the algorithm outputs $(\mathbf{u},\mathbf{v})$.
(2) If $t=2k$, randomly select a non-zero vector $\mathbf{u}\in\mathcal{X}$ and set $\mathbf{v}=\mathbf{w}+\mathbf{u}$. Then $\mathbf{w}=\mathbf{u}+\mathbf{v}$ and $\|\mathbf{v}\|=2k-(k-1)=k+1$, so $\mathbf{v}\in\mathcal{Z}$. The algorithm outputs $(\mathbf{u},\mathbf{v})$.
(3) If t=k1t=k-1, set $\mathbf{u}=\mathbf{w}$ and $\mathbf{v}=0$. This satisfies the output conditions, so the algorithm outputs $(\mathbf{u},\mathbf{v})$.
(4) If $t=k+1$, set $\mathbf{u}=0$ and $\mathbf{v}=\mathbf{w}$. This satisfies the output conditions, so the algorithm outputs $(\mathbf{u},\mathbf{v})$.
(5) If $2\leq t<k-1$, suppose the coordinates of $\mathbf{w}$ equal to 1 are $i_{1},i_{2},\dots,i_{t}$.
  • $\bullet$
    Compute a pair $(n,m)$ such that $t=n+m$ and nm=2n-m=2;
  • $\bullet$
    Construct $\mathbf{v}$: Select $n$ coordinates from $i_{1},i_{2},\dots,i_{t}$ and set them to 1. Choose k+1nk+1-n coordinates from the remaining 2kt2k-t coordinates and set them to 1. All other coordinates are 0. Denote this vector as $\mathbf{v}$;
  • $\bullet$
    Construct $\mathbf{u}$: Set the remaining $m$ coordinates in $i_{1},i_{2},\dots,i_{t}$ (not selected for $\mathbf{v}$) to 1, and set the $n$ selected coordinates to 0. All other coordinates match $\mathbf{v}$. Denote this vector as $\mathbf{u}$.
Verification: Clearly, $\mathbf{w}=\mathbf{u}+\mathbf{v}$. Moreover, $\|\mathbf{u}\|=m+k+1-n=k+1-(n-m)=k-1$ and $\|\mathbf{v}\|=n+k+1-n=k+1$, so $\mathbf{u}\in\mathcal{X}$ and $\mathbf{v}\in\mathcal{Z}$. The algorithm outputs $(\mathbf{u},\mathbf{v})$.
(6) If $k+1<t\leq 2k-2$, suppose the coordinates of $\mathbf{w}$ equal to 1 are $i_{1},i_{2},\dots,i_{t}$.
  • $\bullet$
    Compute a pair $(n,m)$ such that $t=n+m$ and nm=2n-m=2;
  • $\bullet$
    Construct $\mathbf{v}$: Select $n$ coordinates from $i_{1},i_{2},\dots,i_{t}$ and set them to 1. Choose $\frac{2k-t}{2}$ coordinates from the remaining 2kt2k-t coordinates and set them to 1. All other coordinates are 0. Denote this vector as $\mathbf{v}$;
  • $\bullet$
    Construct $\mathbf{u}$: Set the remaining $m$ coordinates in $i_{1},i_{2},\dots,i_{t}$ (not selected for $\mathbf{v}$) to 1, and set the $n$ selected coordinates to 0. All other coordinates match $\mathbf{v}$. Denote this vector as $\mathbf{u}$.
Verification: Clearly, $\mathbf{w}=\mathbf{u}+\mathbf{v}$. Moreover,
$\|\mathbf{u}\|$ $=m+\frac{2k-t}{2}=m+\frac{2k-(n+m)}{2}=k-1,$
$\|\mathbf{v}\|$ $=n+\frac{2k-t}{2}=n+\frac{2k-(n+m)}{2}=k+1,$
so $\mathbf{u}\in\mathcal{X}$ and $\mathbf{v}\in\mathcal{Z}$. The algorithm outputs $(\mathbf{u},\mathbf{v})$.
In all cases, the algorithm finds a valid decomposition $(\mathbf{u},\mathbf{v})$ in polynomial time. ∎
3.1 Construction
The lattice-based short signature scheme proposed in Reference [20] consists of three polynomial-time algorithms: the key generation algorithm KeyGen, the signing algorithm Sign, and the verification algorithm Verify. Specifically, KeyGen takes the security parameter $n$ as input and outputs a public-private key pair $(pk^{\prime},sk^{\prime})$; Sign takes a message $m$ and the private key $sk^{\prime}$ as input and outputs the corresponding signature $\sigma$; while Verify takes the message $m$, the signature $\sigma$, and the public key $pk^{\prime}$ as input and outputs a binary decision result—outputting $1$ indicates acceptance, and outputting 0 indicates rejection. In the subsequent sections of this paper, we denote this signature scheme as $\mathcal{S}=(\textbf{KeyGen},\textbf{Sign},\textbf{Verify})$. Unless otherwise specified, $\mathcal{S}$ consistently refers to a signature scheme that meets security requirements and possesses an almost-tight security reduction.
In this paper, the message space is $\mathcal{M}=\mathbb{F}_{2}^{k}$, and the signature space is $\Sigma=\mathbb{Z}^{2n}\otimes\mathbb{Z}^{n}$. The definitions of sets $\mathcal{X}$, $\mathcal{Y}$, and $\mathcal{Z}$ are consistent with those in Theorem Theorem 6.
Let $\langle f\rangle=(c_{1},...,c_{k_{0}})$ denote the linear function $f(\mathbf{m}_{1},...,\mathbf{m}_{k_{0}})=c_{1}\mathbf{m}_{1}+...+c_{k_{0}}\mathbf{m}_{k_{0}}$, where $\mathbf{m}_{i}\in\mathbb{F}_{2}^{k}$ and $c_{i}\in\{0,1\}$ for all $i$.
Our linearly homomorphic signature scheme $\mathcal{LS}=(\textbf{HSetup},\textbf{HKeyGen},\textbf{HSign},$ $\textbf{HVerify},\textbf{Evaluate})$ works as follows:
[…]
Let $k=2k_{0}$, $q=\mathrm{poly}(n)$ such that $q\geq(nk)^{2}.$ Let $h=\lfloor\frac{n}{6\log q}\rfloor$, $V=\sqrt{2nk\log q}\cdot\log n$. These parameters, together with the algorithm suite $\mathcal{S}$, are set as the public parameter $pp=(q,k,h,V,\mathcal{S})$ and output, where $\mathcal{S}=(\textbf{KeyGen},\textbf{Sign},\textbf{Verify})$.
-HKeyGen(pppp): Input the public parameters $pp$, then the algorithm outputs a pair of public key and secret key The complete key extraction process is formally defined in Algorithm 1.
Algorithm 1 HKeyGen
1:Public parameters pppp
2:Public key $pk$, Secret key sksk
[…]
4:repeat
5:  Sample $\alpha_{1},\ldots,\alpha_{k}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\mathbb{Z}_{q}^{h}$
6:until $\alpha_{1},\ldots,\alpha_{k}$ are linearly independent
[…]
10:return $(pk,sk)$
[…]
Algorithm 2 HSign
1:Secret key $sk=(T_{A},sk^{{}^{\prime}})$, tag $\tau\in\{0,1\}^{n}$, data message $\mathbf{m}_{i}=(m_{i1},...,m_{ik})\in\mathbb{F}_{2}^{k}$, index $i$
2:Signature $(\sigma_{\tau},\sigma_{i})$
[…]
4:Compute $(\mathbf{u}_{i},\mathbf{v}_{i})\leftarrow\textbf{Decompose}(\mathbf{m}_{i})$ $\triangleright$ $\mathbf{m}_{i}=\mathbf{u}_{i}+\mathbf{v}_{i}$ with $\mathbf{u}_{i}\in\mathcal{X}$, $\mathbf{v}_{i}\in\mathcal{Y}$ or $\mathcal{Z}$
5:if $\mathbf{u}_{i}=\mathbf{0}$ or $\mathbf{v}_{i}=\mathbf{0}$ then
6:  Compute $\mathbf{t}_{i}=\sum_{j=1}^{k}m_{ij}\alpha_{j}$
7:  $\sigma_{i}\leftarrow\textbf{SamplePre}(A,T_{A},\mathbf{t}_{i},V)$
8:else
9:  Compute $\mathbf{t}(\mathbf{u}_{i})=\sum_{j=1}^{k}u_{ij}\alpha_{j}$
10:  Compute $\mathbf{t}(\mathbf{v}_{i})=\sum_{j=1}^{k}v_{ij}\alpha_{j}$
11:  $\sigma(\mathbf{u}_{i})\leftarrow\textbf{SamplePre}(A,T_{A},\mathbf{t}(\mathbf{u}_{i}),V)$
12:  $\sigma(\mathbf{v}_{i})\leftarrow\textbf{SamplePre}(A,T_{A},\mathbf{t}(\mathbf{v}_{i}),V)$
13:  $\sigma_{i}=\sigma(\mathbf{u}_{i})+\sigma(\mathbf{v}_{i})$
14:end if
15:return $(\sigma_{\tau},\sigma_{i})$
[…]
Algorithm 3 Verify
1:Public key $pk=(A,pk^{{}^{\prime}},\alpha_{1},\ldots,\alpha_{k})$, tag $\tau\in\{0,1\}^{n}$, message $\mathbf{m}\in\mathbb{F}_{2}^{k}$, signature $\sigma=(\sigma_{\tau},\sigma_{f})$, linear function $f=(c_{1},\ldots,c_{k_{0}})$ where $c_{i}\in\{0,1\}$
2:1 (valid) or 0 (invalid)
3:Compute $\mathbf{t}=\sum_{i=1}^{k_{0}}c_{i}\mathbf{t}_{i}$ where $\mathbf{t}_{i}=\sum_{j=1}^{k}m_{ij}\alpha_{j}$
4:if $1\leftarrow\textbf{Verify }(pk^{{}^{\prime}},\sigma_{\tau})$, $\|\sigma\|\leq k\cdot V\cdot\sqrt{2kn}$ and $A\cdot\sigma\equiv\mathbf{t}\pmod{q}$ then
5:  return 1 $\triangleright$ Signature is valid
6:else
7:  return 0 $\triangleright$ Signature is invalid
8:end if
[…]
Remark 3: In the vector decomposition of message $\mathbf{m}_{i}$, when one of the components is a zero vector, the proof approach for the related conclusions follows the same methodology as the case where both components are non-zero vectors. Therefore, when demonstrating the correctness, unforgeability, and privacy of the scheme, it suffices to consider the scenario where $\mathbf{m}_{i}$ is decomposed into two non-zero vectors.
The following proves that the above scheme satisfies correctness.
Theorem 7.
The above linearly homomorphic signature scheme satisfies correctness with overwhelming probability.
Proof.
According to the construction of the signature algorithm HSign, the first component of the signature satisfies $\sigma_{\tau}\leftarrow\textbf{Sign}(sk^{\prime},\tau)$. Consequently, it necessarily holds that $\boldsymbol{1}\leftarrow\textbf{Verify}(pk^{\prime},\sigma_{\tau})$. Therefore, in the following, we focus solely on verifying the correctness of the second component of the signature.
(1) Given $(\sigma_{\tau},\sigma_{i})=\operatorname{Sign}(sk,\tau,\mathbf{m}_{i},i)$, let $f=\pi_{i}$ be the projection function, which can be viewed as a special type of linear function, that is, $\pi_{i}(\mathbf{m}_{1},\dots,\mathbf{m}_{k_{0}})=\mathbf{m}_{i}$. In this case, we check the following two verification conditions:
From Theorem 2, it follows that $\|\widetilde{T}_{A}\|\leq\mathcal{O}(\sqrt{h\log q})$ with overwhelming probability. Since $V=\sqrt{2nk\log q}\cdot\log n$, it follows that
$\frac{V}{\|\widetilde{T}_{A}\|}\geq\sqrt{\frac{2nk}{h}}\cdot\log n\geq\sqrt{\log n},$
which means $V\geq\|\widetilde{T}_{A}\|\cdot\omega(\sqrt{\log n})$.
Therefore, by Lemma 5, we have $\|\sigma(\mathbf{u}_{i})\|\leq V\cdot\sqrt{n}$ and $\|\sigma(\mathbf{v}_{i})\|\leq V\cdot\sqrt{n}$ with overwhelming probability. Therefore,
$\|\sigma_{i}\|\leq 2V\cdot\sqrt{n}<k\cdot V\cdot\sqrt{2kn}.$
Since
$A\cdot\sigma_{i}\pmod{q}$ $=A\cdot(\sigma(\mathbf{u}_{i})+\sigma(\mathbf{v}_{i}))\pmod{q}$
$=\mathbf{t}(\mathbf{u}_{i})+\mathbf{t}(\mathbf{v}_{i})$
$=\sum_{j=1}u_{ij}\alpha_{j}+\sum_{j=1}v_{ij}\alpha_{j}$
$=\sum_{j=1}(u_{ij}+v_{ij})\alpha_{j}$
$=\sum_{j=1}m_{ij}\alpha_{j}=\mathbf{t}_{i},$
thus, in the case where $f=\pi_{i}$, the scheme satisfies correctness.
(2) Given $(\mathbf{m}_{1},...,\mathbf{m}_{k_{0}})\in(\mathcal{M})^{k_{0}}$, $\overrightarrow{\sigma}=(\sigma_{\tau 1},...,\sigma_{\tau k_{0}})$, where $\sigma_{\tau i}=(\sigma_{\tau},\sigma_{i})=\textbf{HSign}(sk,\tau,\mathbf{m}_{i},i)$. Let $f=(c_{1},...,c_{k_{0}})$, where $c_{i}\in\{0,1\}$. According to the definition of the algorithm $\textbf{Evaluate}(pk,\tau,f,\vec{\sigma})$, $\sigma_{f}=\sum_{i=1}^{k_{0}}c_{i}\sigma_{i}$ is the signature of $f(\mathbf{m}_{1},...,\mathbf{m}_{k_{0}})$. In this case, we will verify that $\sigma_{f}$ satisfies the following two verification conditions:
Since $\parallel\sigma_{i}\parallel\leq 2V\cdot\sqrt{n}$ holds with overwhelming probability, and $|c_{i}|\leq 1$, we have
$\parallel\sigma_{f}\parallel=\parallel c_{1}\sigma_{1}+...+c_{k_{0}}\sigma_{k_{0}}\parallel\leq k_{0}\cdot\mathrm{max}_{1\leq i\leq k_{0}}\parallel\sigma_{i}\parallel\leq 2k_{0}\cdot V\cdot\sqrt{n}\leq k\cdot V\cdot\sqrt{2kn}.$
Since $(\sigma_{\tau},\sigma_{i})=\textbf{HSign}(sk,\tau,\mathbf{m}_{i},i)$, we have $A\cdot\sigma_{i}\pmod{q}=\mathbf{t}_{i}$. Therefore, $A\cdot c_{i}\sigma_{i}\pmod{q}=c_{i}\mathbf{t}_{i}$. By the linearity of modular arithmetic, we can derive:
$A\cdot\sigma\pmod{q}$ $=A\cdot\left(\sum_{i=1}^{k}c_{i}\sigma_{i}\right)\pmod{q}$
$=\sum_{i=1}^{k}A\cdot c_{i}\sigma_{i}\pmod{q}$
$=\sum_{i=1}^{k}c_{i}\mathbf{t}_{i}=\mathbf{t}.$
To sum up, the scheme satisfies the correctness with overwhelming probability. ∎
4 Security and privacy
In the field of provable security, the security of signature algorithms is usually related to hard computational problems. Typically, such proofs describe a reduction from a certain (assumed) hard computational problem to breaking the security properties defined by the signature system. Suppose there exists an adversary $\mathcal{A}$ who can break a signature scheme with a non-negligible advantage $\epsilon$ within polynomial time $t$. We construct a simulator $\mathcal{C}$ to solve a difficult computational problem with an advantage of $\epsilon^{\prime}$ within time $t^{\prime}$, where
$t^{\prime}=t+T,\quad\epsilon^{\prime}=\frac{\epsilon}{L}.$
If $(t^{\prime},\epsilon^{\prime})\approx(t,\epsilon)$, or more specifically, if $T=o(t)$ and $L$ is a constant or relatively small, then we say that this reduction process is a tight reduction, and we say that the signature scheme is tightly secure under the assumption that the computational problem is hard. If the reduction has a security loss where $L$ is polynomially bounded but not constant, the scheme is said to be almost tightly secure. This requires that the tight security of the signature algorithm be independent of the attacker’s parameters, such as the attacker’s number of queries or its security parameter.
In essence, a tight reduction means that a successful attack algorithm can be transformed into an efficient algorithm for solving difficult computational problems without significantly increasing the running time and or significantly reducing the success probability. Therefore, the tightness of the reduction determines the strength of the security guarantee provided by the security proof: a non-tight reduction provides a weaker security guarantee than a tight reduction. Therefore, it is highly desirable for a signature system to have a tight security reduction.
The signature scheme $\mathcal{S}=(\textbf{KeyGen},\textbf{Sign},\textbf{Verify})$ proposed in [20] achieves almost tight security in the standard model. In what follows, we will prove that, under the standard model, the signature scheme proposed in this work also also achieves almost tight security in the EUF-CMA security model.
Theorem 8.
Suppose $q=\mathrm{poly}(n)$, $k_{0}=\mathrm{poly}(n)$, $k=2k_{0}$, $q\geq(nk)^{2}$, $h=\lfloor\frac{n}{6\log q}\rfloor$, $V=\sqrt{2nk\log q}\cdot\log n$, $\beta=2k\cdot V\cdot\sqrt{2kn}$. If the $\textbf{SIS}_{q,\beta}$ problem is hard, then the above linearly homomorphic signature scheme is almost tightly secure under the EUF-CMA security model.
More specifically, let $\mathcal{A}$ be a polynomial-time adversary that makes at most $q_{s}$ signature queries to the simulator $\mathcal{C}$. If the adversary $\mathcal{A}$ breaks the above signature scheme with advantage $\epsilon$ within time $t$, then the simulator $\mathcal{C}$ can construct a polynomial-time algorithm $\mathcal{B}$ that, within time $t+\mathcal{O}\big(q_{s}T_{\textbf{HSign}}+kT_{\textbf{SampleDom}}+q_{s}T_{\mathcal{S}}\big)$, outputs a solution to the $\textbf{SIS}_{q,\beta}$ problem with nearly the same advantage $\epsilon-\mathcal{S}_{\epsilon}-\mathrm{negl}(n)$.
Here, $T_{\textbf{HSign}}$ represents the time spent in generating a signature, $T_{\textbf{SampleDom}}$ represents the time for one execution of the sampling algorithm SampleDom, $T_{\mathcal{S}}$ represents the running time of the signature scheme $\mathcal{S}$, and $\mathcal{S}_{\epsilon}$ denotes the adversary’s advantage in breaking the signature scheme $\mathcal{S}$.
Proof.
Assume there exists a polynomial-time adversary $\mathcal{A}$ that breaks the aforementioned signature scheme with advantage $\epsilon$ within time $t$. We will now prove that a simulator $\mathcal{C}$ can be constructed, which utilizes the adversary $\mathcal{A}$’s ability to forge signatures to construct an algorithm $\mathcal{B}$ that can output a solution to the $\textbf{SIS}_{q,\beta}$ problem.
Uniformly and randomly select a matrix $A\in\mathbb{Z}_{q}^{h\times n}$ as the problem instance of the SIS problem. It is noted that, with overwhelming probability, matrix $A$ is a row-full-rank matrix. Now the simulator $\mathcal{C}$ generates the public key in the following way:
  1. (1)
    The simulator $\mathcal{C}$ runs the algorithm $\textbf{Setup}(1^{n},k_{0})$ to obtain the public parameter $pp=(q,k,h,V,\mathcal{S})$.
  2. (2)
    Compute $(pk^{{}^{\prime}},sk^{{}^{\prime}})\leftarrow\textbf{KeyGen}$.
  3. (3)
    Sample $k$ vectors using the algorithm $\textbf{SampleDom}(1^{n},s)$, where $s=\frac{V}{\sqrt{k/2}}\geq\omega(\sqrt{\log n})$:
    $\mathbf{x}_{i}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\textbf{SampleDom}(1^{n},s),\quad\text{for}\quad i=1,...,k,$
    and then compute $\alpha_{1}=A\cdot\mathbf{x}_{1}(\mathrm{mod}q),...,\alpha_{k}=A\cdot\mathbf{x}_{k}(\mathrm{mod}q)$.
The simulator $\mathcal{C}$ outputs the public key $pk=(A,pk^{{}^{\prime}},\alpha_{1},...,\alpha_{k})$ to adversary $\mathcal{A}$.
After receiving the public key $pk=(A,pk^{{}^{\prime}},\alpha_{1},...,\alpha_{k})$, the adversary $\mathcal{A}$ selects a sequence of data sets $\mathbf{m}_{1},...,\mathbf{m}_{q_{s}}$ as signature queries, where $\overrightarrow{\mathbf{m}}_{i}=(\mathbf{m}_{i1},...,\mathbf{m}_{ik_{0}})\in(\mathcal{M})^{k_{0}}$. For each $i$ ($i=1,...,q_{s}$), the simulator $\mathcal{C}$ choose $\tau_{i}$ uniformly from $\{0,1\}^{n}$ and gives to $\mathcal{A}$ the tag $\tau_{i}$. The steps for the simulator to return the signature queries to the adversary $\mathcal{A}$ are as follows:
  1. (1)
    Take $\tau_{i}\in\{0,1\}^{n}$, $\mathbf{m}_{ij}\in\{\mathbf{m}_{i1},...,\mathbf{m}_{ik}\}$, where $1\leq i\leq q_{s}$ and $1\leq j\leq k_{0}$.
  2. (2)
    Compute $(\mathbf{u}_{ij},\mathbf{v}_{ij}){\leftarrow}\textbf{Decompose}(\mathbf{m}_{ij})$, where $\mathbf{m_{ij}}=(m_{ij}^{(1)},...,m_{ij}^{(k)})=\mathbf{u}_{ij}+\mathbf{v}_{ij}$, $\mathbf{u}_{ij}=(u_{ij}^{(1)},...,u_{ij}^{(k)})\in\mathcal{X}$, $\mathbf{v}_{ij}=(v_{ij}^{(1)},...,v_{ij}^{(k)})\in\mathcal{Y}$ or $\mathcal{Z}$.
  3. (3)
    Compute $\sigma_{\tau_{i}}\leftarrow\textbf{Sign}(sk^{{}^{\prime}},\tau_{i})$
  4. (4)
    Compute $\sigma(\mathbf{u}_{ij})=\sum_{l=1}^{k}u_{ij}^{(l)}\mathbf{x}_{l}$, and $\sigma(\mathbf{v}_{ij})=sum_{l=1}^{k}v_{ij}^{(l)}\mathbf{x}_{l}$. According to Lemma 8 and Corollary 5, $\sigma(\mathbf{u}_{ij})$ and $\sigma(\mathbf{v}_{ij})$ follow distribution $\mathcal{D}_{\mathbb{Z}^{n},V}$ within a negligible statistical distance.
  5. (5)
    Compute $\sigma_{ij}=\sigma(\mathbf{u}_{ij})+\sigma(\mathbf{v}_{ij}).$
The simulator $\mathcal{C}$ output the signature $(\sigma_{\tau_{i}},\sigma_{ij})$.
Next, we prove that the output of the simulator is the same as the distribution in the real signature scheme (within a negligible statistical distance).
In the real scheme, the public-key matrix $A$ is generated by TrapGen and its distribution is statistically close to the uniform distribution over $\mathbb{Z}_{q}^{h\times n}$. In the simulation, $A$ is sampled uniformly at random from $\mathbb{Z}_{q}^{h\times n}$. Thus, the two distributions are statistically indistinguishable.
Secondly, $\alpha_{1},\alpha_{2},...,\alpha_{k}$ is also indistinguishable from the real scheme. This is because, according to Lemma 4, $\alpha_{j}=A\cdot\mathbf{x}_{j}(\mathrm{\mod}q)$ (j=1,,k)(j=1,...,k) is statistically close to being uniformly distributed.
Finally, we show that the simulated signatures are statistically indistinguishable from the real signatures; crucially, we only need to consider the second component of the signature. Since $\mathbf{x}_{j}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\textbf{SampleDom}(1^{n},s)$ and $s=\frac{V}{\sqrt{k/2}}\geq\omega(\sqrt{\log n})$. Thus, by Lemma 5, $\parallel\mathbf{x}_{j}\parallel\leq\frac{V}{\sqrt{k/2}}\sqrt{n}\leq V\sqrt{n}$ holds with overwhelming probability. Therefore,
$\parallel\sigma(\mathbf{u}_{ij})\parallel=\parallel\sum_{l=1}^{k}u_{ij}^{(l)}\mathbf{x}_{l}\parallel\leq V\cdot\sqrt{2kn}.$
Similarly, $\parallel\sigma(\mathbf{v}_{ij})\parallel\leq V\cdot\sqrt{2kn}$. Thus $\parallel\sigma_{ij}\parallel=\parallel\sigma(\mathbf{u}_{ij})+\sigma(\mathbf{v}_{ij})\parallel\leq 2V\cdot\sqrt{2kn}\leq kV\cdot\sqrt{2kn}$.
In addition:
$A\cdot\sigma(\mathbf{u}_{ij})$ $=A\cdot\sum_{l=1}^{k}u_{ij}^{(l)}\mathbf{x}_{l}$ (2)
$=\sum_{l=1}^{k}u_{ij}^{(l)}A\cdot\mathbf{x}_{l}=\sum_{l=1}^{k}u_{ij}^{(l)}\alpha_{l}=\mathbf{t}(\mathbf{u}_{ij})(\mathrm{mod}q)$
Since $\sigma(\mathbf{u}_{ij})$ statistically follows the distribution $\mathcal{D}_{\mathbb{Z}^{n},V}$, but also satisfies the condition $\sigma(\mathbf{u}_{ij})\in\Lambda_{q}^{\mathbf{t}(\mathbf{u}_{ij})}(A)$. Therefore, according to Lemma 9, $\sigma(\mathbf{u}_{ij})$ statistically follows the distribution $\mathcal{D}_{\Lambda_{q}^{\mathbf{t}(\mathbf{u}_{ij})}(A),V}$. Similarly, $\sigma(\mathbf{v}_{ij})$ statistically follows the distribution $\mathcal{D}_{\Lambda_{q}^{\mathbf{t}(\mathbf{v}_{ij})}(A),V}$.
The adversary can also verify as follows:
$A\cdot\sigma_{ij}$ $=A\cdot(\sigma(\mathbf{u}_{ij})+\sigma(\mathbf{v}_{ij}))=A\cdot\sigma(\mathbf{u}_{ij})+A\cdot\sigma(\mathbf{v}_{ij})$ (3)
$=A\cdot\sum_{l=1}^{k}u_{ij}^{(l)}\mathbf{x}_{l})+A\cdot\sum_{l=1}^{k}v_{ij}^{(l)}\mathbf{x}_{l})$
$=\sum_{l=1}^{k}u_{ij}^{(l)}A\cdot\mathbf{x}_{l}+\sum_{l=1}^{k}v_{ij}^{(l)}A\cdot\mathbf{x}_{l}$
$=\sum_{l=1}^{k}u_{ij}^{(l)}\alpha_{l}+\sum_{l=1}^{k}v_{ij}^{(l)}\alpha_{l}$
$=\sum_{l=1}^{k}(u_{ij}^{(l)}+v_{ij}^{(l)})\alpha_{l}=\mathbf{t}_{ij}.$
Hence, the simulated signature and the real signature are also statistically indistinguishable.
After the adversary $\mathcal{A}$ receives all the queried signatures, it will generate a valid forgery $(\tau^{*},\mathbf{m}^{*},\sigma^{*},f)$, where $\mathbf{m}^{*}=(m_{1}^{*},...,m_{k}^{*})$, $\langle f\rangle=(c_{1},...,c_{k})$ , such that
$\textbf{Verify}(pk,\tau^{*},\mathbf{m}^{*},\sigma^{*},f)=1.$
Since $\mathcal{S}$ is a secure signature scheme, there exists some $i$ such that $\tau^{*}=\tau_{i}$, meaning that the adversary can only perform a II-type forgery. Whether it is a I-type forgery or II-type forgery, the following equation holds:
$A\cdot\sigma^{*}(\mathrm{mod}q)=\sum_{i=1}^{k}m_{i}^{\ast}\alpha_{i}=\sum_{i=1}^{k}m_{i}^{\ast}A\cdot\mathbf{x}_{i}=A\cdot\mathbf{x}^{\ast}(\mathrm{mod}q),$
where $\mathbf{x}^{\ast}=\sum_{i}^{k}m_{i}^{\ast}\mathbf{x}_{i}$. By Lemma 5, $\parallel\mathbf{x}_{i}\parallel\leq s\cdot\sqrt{n}=\frac{V}{\sqrt{k/2}}\cdot\sqrt{n}$ holds with overwhelming probability, so $\parallel\mathbf{x}^{\ast}\parallel\leq V\cdot\sqrt{2kn}\leq k\cdot V\cdot\sqrt{2kn}$. By verification condition (1), $\parallel\sigma^{*}\parallel\leq k\cdot V\cdot\sqrt{2kn}$, thus $\parallel\sigma^{*}-\mathbf{x}^{\ast}\parallel\leq 2k\cdot V\cdot\sqrt{2kn}$.
Let the process of outputting $\sigma^{*}-\mathbf{x}^{\ast}$ be denoted as algorithm $\mathcal{B}$. If $\sigma^{*}-\mathbf{x}^{\ast}\neq 0$, then the simulator outputs $\sigma^{*}-\mathbf{x}^{\ast}$ as a solution to the $\textbf{SIS}_{q,\beta}$, where $\beta=2k\cdot V\cdot\sqrt{2kn}$.
Finally, we discuss the probability of $\sigma^{*}=\mathbf{x}^{\ast}$. According to the properties of the sampling algorithm, that is, the preimages have a conditional min-entropy of $\omega(\log n)$. Therefore
$\mathrm{Pr}[\sigma^{*}=\mathbf{x}^{\ast}]\leq 2^{-\omega(\log n)}\leq\mathrm{negl}(n).$
Considering that the adversary’s advantage in breaking the signature scheme $\mathcal{S}$ is $\mathcal{S}{\epsilon}$ (its security loss is at most polynomial in the security parameter), the advantage of algorithm $\mathcal{B}$ in solving the $\textbf{SIS}{q,\beta}$ problem is $\epsilon-S_{1}-\mathrm{negl}(n)$. ∎
Worst-case to average-case reduction. Based on Theorem 1 from the reference [32], when $q\geq\beta\cdot\omega(\sqrt{n\log n})$ holds, the computational hardness of the $\textbf{SIS}_{q,\beta}$ problem is equivalent to that of approximating the SIVP problem in the worst-case setting with an approximation factor of $\beta\cdot\tilde{O}(\sqrt{n})$. In the Setup algorithm, the requirement $q\geq(nk)^{2}$ serves to ensure that $q$ has a large enough value for Theorem 1 to be applicable.
Tight security. Reference [20] instantiated an efficient pseudorandom function (PRF) based on lattice and number-theoretic assumptions to achieve the first lattice-based short signature scheme $\mathcal{S}$ with almost-tight security. The paper also pointed out that if an efficient, tightly secure PRF based on lattices could be constructed (a significant open problem noted in references [44, 45]), then the signature scheme $\mathcal{S}$ could further achieve tight security. Therefore, the scheme $\mathcal{LS}$ proposed in this work not only satisfies almost-tight security but also holds the potential to be enhanced into a tightly secure scheme in the future.
From the proof of Theorem 8, it can be observed that neither a I-type forgery nor a II-type forgery requires the adversary to specify a target tag $\tau_{g}$ in advance. Therefore, our signature scheme is also almost tightly secure under the U-ST-SCMA security model.
Theorem 9.
Suppose $q=\mathrm{poly}(n)$, $k_{0}=\mathrm{poly}(n)$, $k=2k_{0}$, $q\geq(nk)^{2}$, $h=\lfloor\frac{n}{6\log q}\rfloor$, $V=\sqrt{2nk\log q}\cdot\log n$, $\beta=2k\cdot V\cdot\sqrt{2kn}$. If the $\textbf{SIS}_{q,\beta}$ problem is hard, then the above linearly homomorphic signature scheme is almost tightly secure under the U-ST-SCMA security model.
Finally, we prove that our linearly homomorphic signature scheme is weakly context hiding. We have the following theorem.
Theorem 10.
Assume that $V>k_{0}^{s/2}\cdot\omega(\sqrt{1+\log n})$, then the above linearly homomorphic signature scheme is weakly context hiding.
Proof.
The challenger runs HKeyGen, generating a public-private key pair $(pk,sk)$, and sends it to the adversary. Upon receiving the public-private key pair, assume the adversary outputs $(\mathbf{m}_{0},\mathbf{m}_{1},f_{1},...,f_{s})$ to the challenger, where $\mathbf{m}_{0}=(\mathbf{m}_{01},...,\mathbf{m}_{0k}),\mathbf{m}_{1}=(\mathbf{m}_{11},...,\mathbf{m}_{1k})$ and satisfies $f_{i}(\mathbf{m}_{0})=f_{i}(\mathbf{m}_{1})$, i=1,2,,s.i=1,2,...,s.
For convenience, we treat $\mathbf{m}_{b}$ as a matrix over $\mathbb{F}_{2}^{k_{0}\times k}$, where the $j$-th row corresponds to the $j$-th message in the data set $\mathbf{m}_{b}$. We denote $f_{i}>=(c_{i1},...,c_{ik_{0}})$ and construct the $s\times k_{0}$ matrix $G$, where the $i$-th row is $f_{i}>$.Therefore, we have
$G\mathbf{m}_{0}=G\mathbf{m}_{1}.$
In response, the challenger randomly selects a tag $\tau\in\{0,1\}^{n}$ and a random bit $b\in\{0,1\}$. We define $\sigma_{j}^{(b)}$ as the signature (specifically, its second component) for the $j$-th message in the message list $\mathbf{m}_{b}$, noting that the first component of the signature carries no information about $\mathbf{m}_{b}$. We then construct a $k\times n$ matrix $\Sigma^{(b)}$, where the $j$-th row is $(\sigma_{j}^{(b)})^{\top}$. For each i=1,2,,si=1,2,...,s, let $(\sigma_{\tau},d_{i}^{(b)})=\textbf{Evaluate}(pk,\tau,f_{i},(\sigma_{\tau 1}^{(b)},...,\sigma_{\tau k_{0}}^{(b)}))$. Then, we construct an $s\times n$ matrix $D^{(b)}$, where the $j$-th row is $(d_{j}^{(b)})^{\top}$. According to the definition of Evaluate, we have
$D^{(b)}=G\Sigma^{(b)}.$
For $b=0,1$, $\sigma_{i}^{(b)}$ follows the distribution $\mathcal{D}_{\Lambda^{\tau}+\mathbf{t}_{i}^{(b)},\sqrt{2}V}$, where $\mathbf{t}_{i}^{(b)}=\sum_{j=1}^{k}m_{bi}^{(j)}\alpha_{j}$, and $\mathbf{m}_{bi}=(m_{bi}^{(1)},...,m_{bi}^{(k)})$. By the definition of the signing algorithm, the signatures are independent of each other. Therefore, when $V$ is sufficiently large $(>\mu(G)\eta_{\varepsilon}(\Lambda^{\tau})$, by Theorem 3, the distributions of $D^{(0)}$ and $D^{(1)}$ are statistically close. Therefore, any polynomial-time adversary cannot distinguish between $D^{(0)}$ and $D^{(1)}$ with non-negligible probability and further guess the value of $b$. Now we only need to prove that
$V>k_{0}^{s/2}\cdot\omega(\sqrt{1+\log n})\geq\mu(G)\eta_{\varepsilon}(\Lambda^{\tau}).$
First, since the elements of the matrix $G$ are from the set $\{0,1\}$, then by the Hadamard inequality:
$\mu(G)=\sqrt{det(GG^{\top})}\leq\prod_{i=1}^{s}\parallel\mathbf{g}_{i}\parallel\leq k_{0}^{s/2}<V,$
where $\mathbf{g}_{i}$ represents the $i$-th row of the matrix $G$.
Secondly, According to Lemma 6, there exists a negligible $\epsilon$ such that
$\eta_{\epsilon}(\Lambda^{\tau})\leq\omega(\sqrt{1+\log n})<V$
holds with overwhelming probability. Therefore,
$V>k_{0}^{s/2}\cdot\omega(\sqrt{1+\log n})\geq\mu(G)\eta_{\varepsilon}(\Lambda^{\tau}).$
5 Conclusion
Constructing tightly or almost-tightly secure homomorphic signature schemes over lattices in the standard model poses a fundamental challenge in cryptographic theory. This paper presents the first almost-tightly secure linear homomorphic signature scheme based on the SIS assumption in the standard model, which also achieves adaptive security. Furthermore, our scheme can be upgraded to achieve tight security if an efficient and tightly secure PRF based on lattice hardness assumptions becomes available. Several unresolved challenges persist in current research: As noted in [21], designing more efficient signatures and encryption schemes (such as hierarchical encryption, attribute-based encryption, hierarchical signatures, homomorphic signatures supporting polynomial functions, and hierarchical fully homomorphic signatures) with tight or almost tight security in the standard model remains an open issue.